### v1

# PredIN: Towards Open-Set Gesture Recognition via Prediction Inconsistency

Abstract
Gesture recognition based on surface electromyography (sEMG) has achieved significant progress in human-machine interaction (HMI). However, accurately recognizing predefined gestures within a closed set is still inadequate in practice; a robust open-set system needs to effectively reject unknown gestures while correctly classifying known ones. To handle this challenge, we first report prediction inconsistency discovered for unknown classes due to ensemble diversity, which can significantly facilitate the detection of unknown classes. Based on this insight, we propose an ensemble learning approach, PredIN, to explicitly magnify the prediction inconsistency by enhancing ensemble diversity. Specifically, PredIN maximizes the class feature distribution inconsistency among ensemble members to enhance diversity. Meanwhile, it optimizes inter-class separability within an individual ensemble member to maintain individual performance. Comprehensive experiments on various benchmark datasets demonstrate that the PredIN outperforms state-of-the-art methods by a clear margin. Our proposed method simultaneously achieves accurate closed-set classification for predefined gestures and effective rejection for unknown gestures, exhibiting its efficacy and superiority in open-set gesture recognition based on sEMG.
keywords:
open-set recognition , surface electromyography , ensemble learning , prediction inconsistency , gesture recognition
[…]
1 Introduction
In the human-machine interaction (HMI) paradigm, gesture recognition serves as a foundational task and has been extensively applied across diverse domains [1]. Recently, the development of gesture recognition systems [2, 3, 4] based on surface electromyography (sEMG) signals has been remarkable. However, most of them are confined to closed-set scenarios, where the training and test sets share an identical label space. These closed-set systems lack robustness and reliability in the dynamic and ever-changing real world, which causes them to mistake novel gestures or unintentional motions as known ones and generate false interaction signals. Therefore, a robust gesture recognition system, one which can correctly classify predefined known gestures while identifying unknown gestures in real-world scenarios, is in high demand. Scheirer et al. [5] first described the above demand as open-set recognition (OSR), whose test set contains unknown classes that are not included in the training set.
OSR is an active topic in the field of computer vision, with numerous methods continuously being proposed. However, only a few studies [6, 7] focus on open-set sEMG-based gesture recognition. Due to the inherently random and non-stationary nature of sEMG signals, commonly used methods based on reconstruction or generative models in OSR may not be applicable, particularly in achieving closed-set classification accuracies comparable to discriminative methods [8, 9, 10]. A predominant aspect of existing OSR discriminative methods is to explore the distinctions between known and unknown classes, and design various strategies to enlarge them [11]. Accordingly, a score function is derived based on these distinctions to reject the unknown. A recently popular trend for OSR is employing prototype learning (PL) since it establishes a clear distance distinction between the known and unknown, and demonstrates promising performance [12, 13]. PL methods are able to learn a compact feature space while keeping open space for the unknown.
Beyond the distinction of distance, we reveal that prediction inconsistency within the ensemble learning framework can boost the OSR performance. Within an ensemble learning framework, ensemble members trained with different random initializations can converge to significantly different solutions [14]. This variation causes the ensemble model to perform better than any individual members and more diversity among ensemble members leads to better performance, as they typically do not make the same errors on the same inputs [14], which fulfills one of the OSR tasks’ objectives, a satisfactory closed-set classification ability. Unexpectedly, the ensemble diversity also plays a crucial role in identifying the unknown according to our findings, which accomplishes the other objective of OSR tasks well. In this paper, we first discover prediction inconsistency for unknown samples within the ensemble learning framework. Specifically, ensemble members tend to exhibit inconsistent predictions for the unknown (Fig. 1(b)), while consistently agreeing on the same correct results for known ones (Fig. 1(a)). To better understand the prediction inconsistency, it is important to note that classification models will assign improperly high confidence for unknown samples and misclassify them into known classes [12]. The distinction in prediction inconsistency facilitates the differentiation of unknown samples. In our example of the standard ensemble model, which combines two identical networks, diversity between two members is solely attributed to the randomness of the initialization and the learning procedure [15]. Despite this, it is promising that Fig. 1(c) exhibits pronounced distinctions in prediction inconsistency between known and unknown samples. In light of this, a natural idea is to enhance the ensemble diversity in order to magnify the prediction inconsistency for the unknown.
(a) Known samples
(b) Unknown samples
(c) Prediction inconsistency
Fig. 1: An illustration of prediction inconsistency for the unknown between ensemble members. We summarize the prediction results of samples in the two ensemble members and present them as confusion matrices. Each value in the matrices represents the number of known or unknown samples classified into certain known classes within two members. Both horizontal and vertical coordinates represent class labels. There are pronounced distinctions in prediction inconsistency between known and unknown samples. (a) and (b) represent the prediction results of bioDB2 samples. (c) represents the fraction of prediction inconsistency (%) for both known and unknown samples among four public datasets under the standard ensemble model.
To this end, we propose a new ensemble learning approach, PredIN, to magnify the prediction inconsistency by explicitly enhancing ensemble diversity. Specifically, PredIN introduces two complementary losses which regularize the class feature distribution based on prototype learning [12]. Among ensemble members, PredIN maximizes the inconsistency of class feature distribution by inconsistency loss to enhance diversity. Within an individual ensemble member, PredIN incorporates a triplet loss to optimize inter-class separability, thereby maintaining individual performance. PredIN ultimately rejects the unknown based on prediction inconsistency and distance. We conduct comprehensive experiments on public datasets to validate the superiority of our proposed method. The source code is available at https://github.com/Lchenuu/PredIN.
Summary of Contributions
  • 1)
    We reveal the distinction in prediction inconsistency between known and unknown samples within the ensemble learning framework, which significantly facilitates the rejection of unknown classes.
  • 2)
    Based on the above observation, we propose an ensemble learning approach, PredIN, an effective and explicit diversity-inducing method by maximizing the class feature distribution inconsistency, thereby magnifying the distinction in prediction inconsistency.
  • 3)
    Comprehensive experiments on public sEMG datasets demonstrate that our approach simultaneously maintains closed-set classification accuracy for known gestures and improves rejection for unknown gestures, outperforming previous approaches by a clear margin. Additionally, further experiments on public image datasets demonstrate the general applicability of our approach.
The remainder of this paper is structured as follows. In Section 2, we give the related works based on gesture recognition, open-set recognition and ensemble learning. In Section 3, the details of our proposed PredIN are presented. Extensive evaluations compared with state-of-the-art (SOTA) methods and comprehensive analyses of the proposed approach are reported in Section 4 and Section 5. Finally, the conclusion of this paper is presented in Section 6.
2 Related Works
2.1 Closed-set sEMG-based Gesture Recognition
The emergence of deep learning has freed sEMG-based gesture recognition from the constraints of manual feature extraction [16], facilitating a better understanding of human gestures. Various deep learning architectures have been widely employed for this task. Park et al. [17] pioneered the application of Convolutional Neural Network (CNN) models to classify the Ninapro DB2 dataset [18]. Furthermore, more complex CNN models and Recurrent Neural Network (RNN) models have showcased their superiority to fine gesture classification [2, 3]. EMGHandNet [2] proposed a hybrid CNN and Bi-LSTM framework to capture both the inter-channel and temporal features of sEMG. The attention mechanism is also popular in this field, due to the natural electrode channels and spatial attributes of sEMG signals. Sun et al. [19] proposed a multiscale feature extraction network (MSFENet) based on channel-spatial attention to decode the EMG signals.
Although these progresses have been made, closed-set gesture recognition systems are fragile and may generate false interaction signals when facing inferences from novel gestures or intentional muscle contractions, leading to reduced system reliability and user experience. These extraordinary performances of classic closed-set systems are inadequate since their applications are limited when it comes to the real and open world. In contrast, our work aims to develop an open-set gesture recognition system which can correctly classify known gestures while rejecting unknown gestures in real-world scenarios.
2.2 Open-Set Recognition
Open-set recognition seeks to generalize the recognition tasks from a closed-world assumption to an open set. The main challenge that exists in the OSR tasks is the semantic shift where the labels in the training set and testing set are different [20]. Existing methods can be mainly divided into discriminative methods which learn rejection rules directly and generative methods which model the distribution of known or unknown classes [20].
Previous OSR discriminative methods established the rejection rules or distinctions mainly in prediction probability [21, 22] and distance [12, 13, 23]. Bendale and Boult [21] demonstrated the limitation of softmax probabilities and introduced OpenMax, a new model prediction layer based on extreme value theory. CPN [12] was the first to introduce prototype learning to OSR, modeling known classes as prototypes and rejecting the unknown based on distance metric. ARPL [13] considered the potential characteristics of the unknown data and proposed the concept of reciprocal points to introduce unknown information. Subsequently, more methods based on prototype learning have been proposed, focusing on improving the compactness of known features [23], mining high-quality and diverse prototypes [24], or constructing multiple Gaussian prototypes for each class [25]. In addition to distance metric, Park et al. [11] observed the distinction in the Jacobian norm between the known and unknown and devised an m-OvR loss to induce strong inter-class separation within the known classes. Numerous researchers believe that modeling only known classes is insufficient and suggest incorporating prior knowledge about unknown classes by generative models. Some approaches attempted to generate fake data [26], counterfactual images [27] or confused samples [13].
Despite advancing OSR performance in image recognition, only a few studies [6, 7] have focused on the challenge of open-set sEMG-based gesture recognition. Wu et al. [6] identified the unknown based on distinctions in distance and reconstruction error through metric learning and autoencoders (AE). To avoid the high computational complexity of generative models, Wu et al. [7] further introduced the convolutional prototype network (CPN) to construct multiple prototypes for known classes, employing a matching-based approach to reject the unknown. While these methods have made progress, there still needs to be further exploration to enhance the performance of open-set sEMG-based gesture recognition.
Different from the above methods, our approach emphasizes the distinction in prediction inconsistency and distance metric to reject the unknown. In light of these distinctions, we propose a discriminative approach based on the ensemble learning framework, which has been rarely explored in OSR tasks. In this paper, we demonstrate that ensemble learning shows superiority not only in known classification but also in unknown rejection. MEDAF [28] also applied an ensemble learning framework to address OSR, but our motivation differs from theirs as they focus on diverse representations, not prediction inconsistency.
2.3 Ensemble Learning
Ensemble methods benefit from the diversity in predictions among ensemble members, as errors made by some members are mitigated by correct predictions from others [14]. Building on this observation, enhancing ensemble diversity has been a persistent focus in ensemble learning. Considering where diversity is injected among ensemble members, we can categorize existing methods into three main types. The first type focuses on the diversity of the input space. They seek to construct different inputs for each member by bagging [29], data augmentation [30], or orthogonal input gradients [31]. The second type works on the diversity of the weight space with the underlying assumption that an ensemble of neural networks with weights distant from each other produces diversified outputs [32]. For example, Repulsive Deep Ensembles [33] introduced a repulsive term to discourage different ensemble members from collapsing to the same function. The third type focuses on the diversity of the output space including the predictions and features. For instance, DICE [15] increased diversity by reducing spurious correlations among features through a mutual information-based method. DBAT [34] enforced the disagreement of predictions on the auxiliary Out of Distribution (OOD) data to promote diversity.
Similarly, our approach explicitly enhances ensemble diversity by maximizing class feature distribution inconsistency in the feature space. Additionally, our method considers maintaining individual performance, an aspect addressed only in DICE. They aim to achieve an optimal balance between ensemble diversity and individual accuracies.
3 Methodology
3.1 Problem Definition
Considering the sEMG-based gesture recognition in real-world scenarios, we assume that $\mathcal{Y}\subset\mathbb{N}$ is the infinite label space of all possible gesture classes. Assume that $\mathcal{C}=\{1,\ldots,N\}\subset\mathcal{Y}$ represents $N$ known classes of interest. The set $\mathcal{U}=\mathcal{Y}\backslash\mathcal{C}$ represents all unknown classes that need to be rejected. The objective of open-set recognition is to find a measurable recognition function $f^{*}\subset\mathbb{H}$ which minimizes both the empirical classification risk on known samples and the open space risk on unknown samples. Open space risk refers to the risk of incorrectly labeling any unknown class as a known one [5].
[…]
where $D_{c}$ and $D_{u}$ represent samples belonging to known and unknown classes, respectively.
Fig. 2: An illustration of our proposed framework. Our framework ensembles two members, each of which contains an encoder and a set of learnable prototypes. $\mathcal{L}_{PL}$ and $\mathcal{L}_{trip}$ are applied to each member individually while $\mathcal{L}_{incon}$ simultaneously acts on both. $\mathcal{L}_{incon}$ aims to maximize the class feature distribution inconsistency, ensuring each member has an entirely distinct layout or neighboring class pairs (blue arrows). Upon this, unknown samples like $\mathbf{z}_{u}$ represented by purple are predicted near the clusters of different known classes due to prediction inconsistency while known samples agree on the same predictions across two members.
3.2 Methodology Overview
Combined with prototype learning, we propose an ensemble learning approach, PredIN, to minimize both the empirical classification risk and open space risk simultaneously. Our proposed framework is illustrated in Fig. 2. Specifically, we simultaneously train two ensemble members. In the following text, unless otherwise specified, the number of ensemble members is two. Each member contains an encoder $h$ with an arbitrary architecture and $N$ learnable prototypes of known classes. Prototype learning acts as a basic classification model, with the loss $\mathcal{L}_{PL}$ applied to each member to learn a clear distance distinction. Furthermore, to enlarge the distinction in prediction inconsistency, the inconsistency loss $\mathcal{L}_{incon}$ operates concurrently on both members to enhance their diversity by maximizing the inconsistency of class feature distribution among ensemble members. In addition, we apply the loss $\mathcal{L}_{trip}$ to enhance inter-class separability, thereby maintaining the individual performance.
3.3 Prototype Learning
Each individual member of PredIN employs prototype learning to establish a clear distance distinction. The core idea of PL methods is to encourage samples to be close to their corresponding prototypes and distant from others, where class prototypes serve as centers or representatives of each class [12]. This establishes a compact feature space and a closed classification boundary for known classes while preserving open space for unknown samples [13]. It provides a distance-based approach for rejecting unknown samples, which has been proven superior to softmax-based approaches [24].
For a given sample $\mathbf{x}_{i}$ with label $y_{i}$, its embedding feature is defined as $\mathbf{z}_{i}=h(\mathbf{x}_{i})\in\mathbb{R}^{d}$. $N$ known classes are each assigned a learnable prototype $\mathbf{p}^{k}\in\mathbb{R}^{d}$, where $1\leq k\leq N$. The probability of the prediction result $\hat{y_{i}}$ being $k$ for $\mathbf{x}_{i}$ is based on the distance $d(h(\mathbf{x}_{i}),\mathbf{p}^{k})$:
$p(\hat{y_{i}}=k|\mathbf{x}_{i},h,\mathbf{p})=\frac{e^{-d(h(\mathbf{x}_{i}),\mathbf{p}^{k})}}{\sum_{j=1}^{N}e^{-d(h(\mathbf{x}_{i}),\mathbf{p}^{j})}}.$ (2)
To narrow the distance between samples and their corresponding prototypes while pushing them away from other prototypes, the DCE loss function [12] is utilized and described as follows:
$\mathcal{L}_{\epsilon}=-\frac{1}{M}\sum_{i=1}^{M}\log{(p(\hat{y_{i}}=k|\mathbf{x}_{i},h,\mathbf{p}))},$ (3)
where $M$ represents the number of known samples. We use the dot product to measure the generalized distance between the samples and prototypes as
$d(\mathbf{z}_{i},\mathbf{p}^{k})=-{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}.$ (4)
The DCE loss only guarantees the discriminability of feature space. To enhance the intra-class compactness, we incorporate an additional compactness term as
$\mathcal{L}_{com}=\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{n}(h(\mathbf{x}_{i})-\mathbf{p}^{k}),$ (5)
in which
$\mathcal{L}_{n}(\mathbf{u})=\begin{cases}\frac{1}{2}\|\mathbf{u}\|_{2}&\quad\|\mathbf{u}\|_{1}<1\\ \|\mathbf{u}\|_{1}-\frac{1}{2}&\quad\|\mathbf{u}\|_{1}\geq 1\\ \end{cases},$ (6)
where $y_{i}=k$.
Combining Eq. (3) and Eq. (5), the overall loss function $\mathcal{L}_{PL}$ of the PL baseline is expressed as follows:
$\mathcal{L}_{PL}=\mathcal{L}_{\epsilon}+\beta\mathcal{L}_{com},$ (7)
where $\beta$ controls the intensity of $\mathcal{L}_{com}$.
3.4 Ensemble Class Feature Distribution Inconsistency
In deep learning, class feature distribution is derived from the projection of high-dimensional sample space to low-dimensional feature space with an encoder. Given this, maximizing class feature distribution inconsistency among ensemble members can promote diverse mapping functions, thereby enhancing ensemble diversity. From a global perspective, class feature distribution inconsistency refers to the entirely distinct layout or relative positions of deep features from different classes. Locally, it means that the neighboring classes for each class among ensemble members differ, as illustrated in Fig. 3.
Fig. 3: An illustration of class feature distribution inconsistency between ensemble members. Different colors represent different classes.
Based on PL, the class feature distribution can be approximately characterized by learned prototypes as the actual distribution is not directly accessible and prototypes provide the first-order statistics of the distribution [35]. We compute the distance between each sample feature $\mathbf{z_{i}}$ and prototypes $\mathbf{p}^{j}$ of non-corresponding classes from the same ensemble member:
$d(\mathbf{z}_{i},\mathbf{p}^{j})=-{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{j}},$ (8)
where $1\leq j\leq N$ and $j\neq y_{i}$.
To compare class feature distribution across different feature spaces, we convert distances into probabilities that signify proximity. In this procedure, we have two criteria: firstly, the closer the distances between features, the higher the corresponding probability values. Secondly, adjusting the relative positions between two widely separated classes has a relatively minor impact on the change of class feature distribution compared to adjusting local structures. Therefore, we aim to focus more on neighboring classes. Softmax meets these two requirements, which adjusts the global layout but highlights the influence of neighboring classes:
$p(\mathbf{z}_{i},\mathbf{p}^{k})=\frac{e^{-d(\mathbf{z}_{i},\mathbf{p}^{k})}}{\sum_{j=1}^{N}e^{-d(\mathbf{z}_{i},\mathbf{p}^{j})}},$ (9)
where $1\leq k,j\leq N$ and $k,j\neq y_{i}$.
The collection of probabilities represents the class feature distribution within each member, depicted as an $M\times(N-1)$ matrix in Fig. 2. The computation of class feature distribution involves both sample features and prototypes, ensuring that every sample in the feature space contributes to the adjustment in class feature distribution, not just the representatives (prototypes) of feature clusters. To maximize class feature distribution inconsistency, we propose an inconsistency loss which encourages maximizing the inconsistency of two probabilities as follows:
$\mathcal{L}_{incon}=-\frac{1}{M}\sum_{i=1}^{M}\log{\sum_{k=1}^{N}({p_{m_{A},i}^{k}}(1-{p_{m_{B},i}^{k}})+{p_{m_{B},i}^{k}}(1-{p_{m_{A},i}^{k}}))},$ (10)
where $k\neq y_{i}$. Here $m_{A}$ and $m_{B}$ represent two ensemble members, and $p_{m_{A},i}^{k}$ and $p_{m_{B},i}^{k}$ represent the respective probability values in the Eq. (9). From a mathematical perspective, the inconsistency loss is minimized when two probability distributions take opposite extreme values. In terms of class feature distribution, minimizing inconsistency loss $\mathcal{L}_{incon}$ narrows the proximity between two classes within one member while simultaneously increasing the proximity between the corresponding two classes in another member, as shown in Fig. 4.
Fig. 4: An illustration of how inconsistency loss $\mathcal{L}_{incon}$ acts on two members to adjust the class feature distribution. The class pairs of each member are optimized in opposite directions. When the class pairs are pulled within the margin, the optimization of $\mathcal{L}_{incon}$ will halt for one member.
Considering that adjusting the class feature distribution will inevitably pull some features and non-corresponding prototypes close within a member during training, we introduce the positive distance between features and their corresponding prototypes along with a margin $m_{1}>0$ in the distance computation of the Eq. (8) to mitigate this issue:
[…]
where $1\leq j\leq N$ and $y_{i}=k\neq j$. The term ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}$ is not subject to gradient optimization. The redefined distance metric not only considers the relationship with positive pairs ${\mathbf{z}_{i}}$ and ${\mathbf{p}^{k}}$ but also ensures that the inter-class distance does not decrease to within the margin. Specifically, when ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{j}}>{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}-m_{1}$, the relative positions of corresponding features and prototypes will not be adjusted in one member, as shown in Fig. 4.
3.5 Individual Inter-class Separability
Ensemble learning strategies encourage diversity among members and therefore increase their bias, which may potentially degrade individual performance [15]. While enhancing diversity among ensemble members, we also focus on maintaining individual performance within each member. During maximizing class feature distribution inconsistency, our redefined distance metric ensures a margin for neighboring classes but does not push them away. Since individual ensemble members rely on distance metrics for rejection, establishing a decision boundary with effective inter-class separability is crucial. We therefore introduce the triplet loss [36] based on prototype learning to optimize the inter-class separability. Triplet loss minimizes the distance between an anchor and a positive, both of which belong to the same class, and minimizes the distance between the anchor and a negative of a different class [36]. Neighboring classes naturally form hard negative pairs.
[…]
where $k=y_{i}$. Class $j$ is the nearest neighbor of class $k$. The loss $\mathcal{L}_{trip}$ pulls the features (anchors) and their corresponding prototypes (positives) closer while pushing them away from the nearest non-corresponding prototypes (negatives), ensuring effective inter-class separability for each individual member.
The final loss function applied to PredIN is as follows:
$\mathcal{L}_{Div}=\mathcal{L}_{PL}+\gamma\mathcal{L}_{incon}+\alpha\mathcal{L}_{trip},$ (13)
where $\gamma$ and $\alpha$ are the weights of $\mathcal{L}_{incon}$ and $\mathcal{L}_{trip}$. Here the inconsistency loss and triplet loss are complementary as the former enhances ensemble diversity by maximizing the class feature inconsistency among ensemble members while the latter maintains the individual performance by optimizing the inter-class separability within each member.
3.6 Unknown Rejection
In PredIN, each ensemble member follows previous PL-based approaches and obtains similarity based on the distance between sample features and prototypes. Specifically, similarity between a given feature $\mathbf{z}_{i}$ and the prototype $\mathbf{p}^{k}$ for each member is defined as follows:
$\textrm{Sim}(\mathbf{z}_{i},\mathbf{p}^{k})={\mathbf{z}_{i}}^{T}\cdot\mathbf{p}^{k},$ (14)
where $1\leq k\leq N$.
Fig. 5: Rejection Rules. Through ensemble averaging, unknown samples (right) tend to predict different results and produce lower $S_{max}$, while known samples (left) tend to obtain the same predictions consistent with the label (yellow) and produce higher $S_{max}$.
Besides the above distance metric, we also need to address how to leverage prediction inconsistency brought by ensemble diversity. In ensemble learning, a commonly used method for integrating predictions from different members is averaging [14, 37]. This approach is equally applicable to our task. We combine the outputs of ensemble members:
$\textrm{Score}(\mathbf{z}_{i},\mathbf{p}^{k})=\frac{1}{2}({\textrm{Sim}_{A}(\mathbf{z}_{i},\mathbf{p}^{k}})+{\textrm{Sim}_{B}(\mathbf{z}_{i},\mathbf{p}^{k})}),$ (15)
where $1\leq k\leq N$.
In order to classify sample $\mathbf{x}_{i}$ as a certain known class $k$ or reject it as unknown classes, we further obtain $S_{max}$:
$S_{max}=\textrm{Score}(\mathbf{z}_{i},\mathbf{p}^{k^{*}}),$ (16)
where
$k^{*}=\underset{1\leq k\leq N}{\arg\max}\ \textrm{Score}(\mathbf{z}_{i},\mathbf{p}^{k}).$ (17)
In summary, we comprehensively consider the distinctions in prediction inconsistency and distance between known and unknown classes, thereby effectively increasing the rejection performance of the unknown. As shown in Fig. 5, in a single member, the maximum similarity score of some unknown samples may be higher than that of known samples, making it difficult to reject based on a certain threshold. However, unknown samples will obtain lower $S_{max}$ through ensemble averaging due to prediction inconsistency, while known samples will be close to the same-class prototypes among ensemble members resulting in higher $S_{max}$. A pre-determined threshold can be applied to $S_{max}$ to reject the unknown. Samples with $S_{max}$ greater than the threshold value will be regarded as known ones and classified into class $k^{*}$.
4 Experiments and results
4.1 Datasets
We apply four public sEMG benchmark datasets [38, 18, 39, 40] to validate the proposed approach, as shown in Table 1. During preprocessing, raw sEMG signals are segmented via a sliding window of length 200 ms with steps of 50 ms, and then standardized channel-wise. As recommended by BioPatRec [38], we remove transient periods of the contraction using a contraction time percentage of $0.7$ for the BioPat DB2 dataset. According to the setting of closed-set gesture recognition based on sEMG [2, 3, 4], the training and testing set are divided based on trials as mentioned in Table 1. Following the protocol of open-set image recognition [13], we randomly select $10$ known classes from BioPat DB2 and $15$ known classes from Ninapro DB2, Ninapro DB4 and Ninapro DB7, while treating the remaining classes as unknown.
Table 1: Characteristics and setup of four public sEMG datasets.
Dataset BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
Subjects 17 40 10 20
Channels 8 12 12 12
Sampling rate 2000Hz 2000Hz 2000Hz 2000Hz
Trials 3 6 6 6
Training Trials 1,2 1,3,4,6 1,3,4,6 1,3,4,6
Testing Trials 3 2,5 2,5 2,5
Gestures 27 50 53 41
Known Gestures 10 15 15 15
Unknown Gestures 17 35 38 26
4.2 Evaluation Metrics
We use three common metrics to measure the performance of OSR derived from [13, 11]: (1) the area under the receiver operating characteristic (AUC); (2) closed-set classification accuracy (ACC); (3) open-set classification rate (OSCR). They are all threshold-independent metrics. Further details are provided as follows:
  • $\bullet$
    AUC measures the model’s ability to distinguish between known and unknown classes based on the relationship between true positive rate (TPR) and false positive rate (FPR).
  • $\bullet$
    ACC assesses known classes classification performance.
  • $\bullet$
    OSCR comprehensively evaluates empirical classification risk and open space risk based on closed-set classification accuracy (ACC) and false positive rate (FPR).
4.3 Experimental Settings
To comprehensively extract sEMG features, we employ two types of encoders, the Crossformer [41] and a CNN-LSTM hybrid network, as the backbone. The Crossformer is a popular Transformer-based model for time series forecasting while also showing superiority in sEMG classification tasks since it effectively captures the cross-time and cross-channel dependency and extracts multi-scale time information. We set the segment length of Crossformer as $32$, window size as $2$, and layers as $5$. In addition, we design a hybrid network based on CNN and LSTM, which combines a ResNet variant [42], an LSTM [43] and an SKAttention module [44].
During experiments, we use the SGD optimizer with an initial learning rate of $0.01$. The learning rate decreases by a factor of $0.1$ at $60$ and $80$ epochs. The batch size is set to $256$ and the training epoch is set to $100$. The hyperparameters $\beta$, $\gamma$ and $\alpha$ in Eq. (7) and Eq. (13) are all empirically set to $1.0$, while the feature dimension of embedding space is set to $128$. Two margins in Eq. (11) and Eq. (12) are set to $0.5$ and $1.0$ respectively. Prototypes are randomly initialized by the standard normal distribution. All experimental results are averaged among five randomized splits of datasets by classes, which means each split uses different known classes to train. The PredIN is implemented by using Pytorch 2.3.0 and executed on an NVIDIA GeForce RTX 4090 GPU.
Table 2: Performance comparison with SOTAs in terms of AUC (%) and OSCR (%) on four public datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold.
Methods BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
 AUC OSCR  AUC OSCR  AUC OSCR  AUC OSCR
Softmax 65.5 60.5 72.2 65.6 73.1 61.7 73.5 66.8
ARPL (TPAMI’22) 68.2 61.1 76.3 67.9 79.6 64.5 79.1 70.1
SLCPL (CVIU’23) 70.1 63.7 76.5 68.4 76.4 63.3 76.6 68.4
DIAS (ECCV’22) 70.2 65.8 77.8 69.7 80.2 65.7 79.8 71.0
CPN-MGR (IEEE Sens. J’22) 69.7 63.3 76.4 68.5 76.6 63.2 76.4 68.4
PredIN (Crossformer) 74.9 69.8 80.4 72.5 80.8 68.2 81.0 73.1
OpenMax (CVPR’16) 68.6 64.7 69.2 58.4 74.5 61.2 71.1 62.2
ARPL (TPAMI’22) 73.4 70.0 72.2 59.1 79.8 63.0 78.6 64.6
MGPL (Inf. Sci’23) 70.6 67.3 62.3 44.7 76.7 58.3 64.9 51.7
MEDAF (AAAI’24) 74.7 72.2 72.9 63.2 80.2 64.9 79.9 70.5
CPN-MGR (IEEE Sens. J’22) 73.0 69.3 71.9 55.4 79.3 58.0 78.2 63.6
PredIN (Hybrid model) 75.7 72.7 76.7 64.3 82.2 67.6 82.1 71.6
Table 3: Performance comparison with SOTAs in terms of closed-set ACC (%) on four public datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold.
Methods  BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
Softmax 86.0 82.6 75.3 83.0
ARPL (TPAMI’22) 85.7 82.7 74.8 82.9
SLCPL (CVIU’23) 85.8 82.5 75.5 82.8
DIAS (ECCV’22) 88.1 83.5 75.9 83.5
CPN-MGR (IEEE Sens. J’22) 85.9 82.6 75.0 83.1
PredIN (Crossformer) 89.2 85.2 78.1 85.2
OpenMax (CVPR’16) 87.8 74.7 73.0 78.2
ARPL (TPAMI’22) 91.9 74.6 73.1 78.5
MGPL (Inf. Sci’23) 91.5 65.4 70.1 68.9
MEDAF (AAAI’24) 93.0 77.7 74.4 82.4
CPN-MGR (IEEE Sens. J’22) 92.0 69.5 67.7 75.3
PredIN (Hybrid model) 93.4 77.6 76.9 82.2
4.4 Comparison with the State-of-the-arts
We compare our method against other state-of-the-art open-set image and gesture recognition approaches. Softmax and OpenMax are methods based on the prediction probability. OpenMax [21] uses extreme value theory (EVT) to calibrate the prediction probability. ARPL [13] and SLCPL [23] are methods based on prototype learning and design various distance loss functions to reduce open space risk. DIAS [26] considers different difficulty levels and introduces an image generator and a feature generator to produce hard fake instances. MGPL [25] is also the method based on PL but applies the VAE framework to optimize generative constraints. MEDAF [28] is the method based on ensemble learning, which encourages multiple experts to learn diverse representation with an attention diversity regularization. CPN-MPR [7] focuses on open-set sEMG-based gesture recognition and introduces the PL to reject the unknown. To ensure the fairness of the comparison, all methods employ the same backbone.
The results in Table 2 and Table 3 highlight the performance of our proposed approach for the open-set sEMG-based gesture recognition task, even though across different backbone architectures. Specifically, considering the rejection performance, our method achieves the best AUC scores of 75.7%, 76.7%, 82.2% and 82.1% across four datasets. Moreover, ensemble learning brings benefits to closed-set classification tasks. Our method achieves significant improvements in closed-set accuracy compared to these SOTA methods. Furthermore, when considering both empirical classification risk and open space risk, our approach also surpasses the above SOTA methods, consistently achieving the highest OSCR scores on four datasets. These results confirm that prediction inconsistency reveals the distinctions between known and unknown classes effectively. In conclusion, our approach shows the superiority in both closed-set classification and unknown rejection.
Table 4: Ablations of each module in terms of AUC (%) and OSCR (%) on four public datasets. Best performances are highlighted in bold.
Methods BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
 AUC OSCR  AUC OSCR  AUC OSCR  AUC OSCR
PL baseline 71.5 66.5 73.3 59.1 79.1 61.7 77.7 65.4
Deep Ensemble 73.2 68.9 76.0 64.0 80.9 65.8 80.2 70.2
Deep Ensemble (w/ $\mathcal{L}_{trip}$) 73.8 70.7 76.4 64.2 81.8 67.4 80.8 71.0
PredIN (w/o $\mathcal{L}_{trip}$) 72.4 68.3 75.6 61.9 81.6 66.4 81.1 70.6
PredIN 75.7 72.7 76.7 64.3 82.2 67.6 82.1 71.6
4.5 Ablation Study
Module Ablation. As presented in Table 4, each component within our method has been systematically integrated into the PL baseline to verify its necessity. We first introduce the Deep Ensemble framework, which combines two identical networks with different random initializations. As shown in the second raw, the introduction of ensemble learning yields notable improvements in AUC and OSCR scores compared to the PL baseline, demonstrating the effectiveness of using prediction inconsistency for unknown rejection. Secondly, we further verify the effectiveness of our proposed two losses. In PredIN, the roles of inconsistency loss and triplet loss are complementary: the loss $\mathcal{L}_{trip}$ acts to optimize the inter-class separability with each ensemble member while the inconsistency loss $\mathcal{L}_{incon}$ maximizes the class feature distribution inconsistency among members. To further explain, we apply the loss $\mathcal{L}_{trip}$ to the Deep Ensemble model alone and then combine these two losses into the model training. The standalone application of $\mathcal{L}_{trip}$ enhances the individual rejection capability, thereby consistently improving overall ensemble performance compared to the Deep Ensemble model. When combining these two losses, the PredIN magnifies the prediction inconsistency by enhancing the ensemble diversity while maintaining the individual performance, which provides further improvements and clearly demonstrates the effectiveness of maximizing the class feature distribution inconsistency. In addition, we remove the loss $\mathcal{L}_{trip}$ from the PredIN to verify the importance of maintaining individual performance from another perspective. A clear decrease occurs on BioPat DB2 and Ninapro DB2 after the removal. Finally, incorporating all the above components improves AUC by an average of +3.8% compared to the baseline PL model, which demonstrates that each component contributes to the overall improvement on unknown rejection.
Hyperparameters Ablation. We evaluate the effect of two sets of hyperparameters: the trade-off weights $\beta$ and $\gamma$ in Eq. (7) and Eq. (13), and two margins $m_{1}$ and $m_{2}$ in Eq. (11) and Eq. (12). These experiments are conducted using the bioDB2 dataset. Fig. LABEL:fig6:a shows the AUC for different $\beta$ and $\gamma$ values. The weight $\beta$ influences the compactness of feature space in prototype learning. According to the results, larger values yield better results, but excessively large values may cause optimization issues and lead to a performance decrease. The weight $\gamma$ represents the degree of adjustment to the class feature distribution. Larger values similarly affect the model’s classification. Setting $\beta$ and $\gamma$ to $1.0$ is optimal. Two margins influence the degree of class feature distribution inconsistency and inter-class separability. Greater values of $m_{1}$ and $m_{2}$ lead to reduced inconsistency but increased enhancement for inter-class separability. A balanced result is that the optimal values for $m_{1}$ and $m_{2}$ are set to $0.5$ and $1.0$ in our experiments, respectively.
5 Further Analysis and Discussion
5.1 Image Domain Verification
Table 5: Performance comparison (%) with SOTAs in terms of AUC on four public image datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold. * indicates the reproduced result to unify the split information.
Methods MNIST SVHN CIFAR10 TinyIN
Softmax 97.8 88.6 67.7 57.7
OpenMax (CVPR’16) 98.1 89.4 69.5 57.6
CROSR (CVPR’19) 99.1 89.9 88.3 58.9
PROSER (CVPR’21) - 94.3 89.1 69.3
CPN (TPAMI’22) 99.0 92.6 82.8 63.9
ARPL (TPAMI’22) 99.6 96.3 90.1 76.2
ODL (TPAMI’22) 99.6 95.4 88.5 74.6
SLCPL (CVIU’23) 99.4 95.2 86.1 74.9
MGPL (Inf.Sci’23) - 95.7 84.0 73.0
m-OvR (Pattern Recognit.’24) - 95.7 89.5 75.3
DIAS* (ECCV’22) 99.5 94.7 90.3 76.8
PredIN 99.6 97.2 90.5 77.2
To further verify the performance of our proposed approach, we make a comparative experiment on four public image datasets widely used for OSR performance evaluation, including MNIST, SVHN [45], CIFAR10 [46] and TinyImageNet [47]. We compare our proposed method, PredIN, to the SOTA open-set recognition methods including Softmax, OpenMax [21], CROSR [48], PROSER [22], CPN [12], ARPL [13], ODL [49], SLCPL [23], MGPL [25], m-OvR [11] and DIAS [26]. For a fair comparison, we use the same backbone VGG [50] as these methods. In terms of optimization, we use the SGD optimizer with a momentum value of $0.9$ and set the initial learning rate to $0.01$ which drops to $0.1$ at every 30 epochs. The parameters $\beta$, $\gamma$, $\alpha$, $m_{1}$, $m_{2}$ and feature dimension are set to $1.0$, $1.0$, $1.0$, $0.5$, $0.5$ and $128$, respectively. All results of SOTA methods are taken from the references except DIAS [26]. As DIAS [26] applies different dataset split ways, we reproduce their method using the recommended hyperparameters to unify the split information. AUC performances are shown in Table 5. Our approach achieves comparable performance to SOTA methods. This clearly demonstrates the general applicability of our approach across sEMG and image domains.
(a) Deep Ensemble model
(b) PredIN
Fig. 7: The visualization of class feature distribution of two ensemble members. Both horizontal and vertical coordinates represent class labels. Each value represents the degree of proximity between two classes within a single member. (a) represents the two members of the Deep Ensemble model. There are similar local structures between the two members (yellow boxes). (b) represents the two members of the PredIN model. The global layout and local neighboring pairs are different (red boxes).
5.2 Ensemble Diversity Analysis
The improvement in performance compared to the Deep Ensemble model confirms the increased ensemble diversity introduced by PredIN. To further evaluate the diversity, we visualize the class feature distribution and measure the diversity with a diversity metric $div$.
We first verify that our approach achieves class feature distribution inconsistency. Specifically, we visualize the class feature distribution using a proximity matrix. As the class feature distribution can be approximately characterized by learned prototypes, we compute the distances between class prototypes and convert them into probabilities in order to represent class proximity. Fig. 7 demonstrates that PredIN achieves our desired class feature distribution inconsistency between ensemble members, especially compared to the Deep Ensemble model. Locally, each class has different neighboring classes, while globally, the relative positions of classes vary, which causes the feature spaces of the two members different.
Fig. 8: Diversity comparison on four sEMG datasets.
To further validate that maximizing the class feature distribution inconsistency enhances ensemble diversity, we compare the PredIN to the Deep Ensemble model over a diversity metric. Common diversity metrics [14, 15] employed in classification tasks focus on the fraction of label changes in known samples and are therefore not directly applicable to OSR. Consequently, we modify these metrics to measure the relative prediction inconsistency between ensemble members on unknown samples:
$div=\frac{Fraction\ of\ unknown\ label\ changes}{Fraction\ of\ known\ label\ changes}.$ (18)
The results in Fig. 8 demonstrate the consistent diversity improvements achieved by our approach across four sEMG datasets.
5.3 Ensemble Number
In our experiments, we use two ensemble members because the inconsistency loss has a symmetrical form, which makes it unsuitable for optimizing more members in parallel within the ensemble. To evaluate whether more ensemble members will perform better, we train multiple models sequentially. The training of the first model is only based on the PL loss $\mathcal{L}_{PL}$, minimizing its empirical classification risk. Subsequent models apply our proposed approach and are optimized based on the loss $\mathcal{L}_{Div}$. Each model obtains a different class feature distribution from the former. Fig. LABEL:fig10 shows the results for a larger ensemble number of $5$ on bioDB2 and DB7. The more ensemble models gain better performance, which aligns with the assumption of ensemble learning. However, ensembling more members will lead to performance saturation and increased computational burden.
5.4 Limitations and Future Work
Although our proposed PredIN addresses the challenge of open-set gesture recognition and demonstrates promising performance, it has several limitations. One limitation is the computation cost associated with the ensemble learning framework. Ensemble learning brings diversity but also introduces certain computational costs. In future work, we can mitigate this by sharing a shallow encoder among ensemble members. Additionally, enhancing the diversity of the ensemble and improving the performance of individual members are orthogonal objectives. By developing an individual member with optimal performance and combining it with the diversity of the ensemble model, we can further enhance the model’s rejection performance. The potential of ensemble models in OSR tasks remains to be fully explored. Moreover, our approach is limited to the rejection of unknown gestures, without further considering the system’s ability to learn and classify novel gestures dynamically. One potential future work is to incorporate open-world recognition frameworks, such as leveraging incremental learning techniques.
6 Conclusion
Generalizing gesture recognition from closed-set to open-set is important for real-world HMI. To tackle open-set gesture recognition based on sEMG, we propose an ensemble learning approach, PredIN, based on our observed prediction inconsistency for the unknown due to the ensemble diversity. Specifically, we propose two complementary losses to improve OSR performance by enhancing ensemble diversity while maintaining individual performance. Extensive experiments conducted on multiple datasets consistently demonstrate that our approach outperforms previous state-of-the-art open-set classifiers. This means that our gesture recognition system can maintain high classification accuracy for predefined gestures, while effectively rejecting gestures of disinterest. We hope this work will boost the applications of gesture recognition technologies in real-world scenarios. Moving forward, we will also explore extending our technology to adapt to diverse recognition tasks.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgments
This work was partially supported by OYMotion Technologies.